OpenStudy (anonymous):

Solve the following equation analytically. (Enter your answers as a comma-separated list.) log27(3x − 1)/ (2x + 4) = 1/3

OpenStudy (anonymous):

Can you clarify the equation? It's a little ambiguous as it sits. Is 27 the base of the log? Is the 2x+4 outside the log or inside?

hartnn (hartnn):

is it \(\log_{27}[\dfrac{(3x − 1)}{ (2x + 4)}] = 1/3\) ??

OpenStudy (anonymous):

yea

hartnn (hartnn):

you can use the log property that \(\large \log_ab=c \implies a^c=b\) what will u get for your question ?